# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

### v1

@@ 4.1 Experimental Settings
We verify our ViSurf on the following benchmarks across several domains.
Non-Object Segmentation. The gRefCOCO [16] includes queries that do not contain corresponding objects. The evaluation metrics are gIoU and N-Acc. We use Multi-objects-7K plus with 200 non-object data for training.
Reasoning Segmentation. The ReasonSeg [12] includes test samples that need reasoning for correct segmentation. It has 200 validation images and 779 test images. The evaluation metric is gIoU. We use Multi-objects-7K proposed in VisionReasoner [21] for training.
GUI Grounding. The OmniACT [11] is a GUI grounding task for Desktop and Web. We derive 6,101 samples in training split and verify on the test split. The accuracy is calculated as whether the predict point correctly locates in the interest region.
Anomaly Detection. The RealIAD [35] includes real-world, multi-view industrial anomaly. We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint. We calculated the ROC_AUC.
Medical Image: Skin. The task one of ISIC2018 [4, 10] is lesion segmentation. It includes 2,594 training samples and 1,000 test samples. We measure the bbox_acc metric, which computes the ratio of predicted bounding boxes whose IoU with the ground truth exceeds 0.5.
MathVista. The MathVista-testmini [24] includes 1,000 diverse mathematical and visual tasks. We gather around 10k training data from WeMath [29], MathVision [37], Polymath [8], SceMQA [15], Geometry3K [23].
Implemention Details. We choose Qwen2.5-VL [1] and SAM2 [31] (if applicable) as the baseline model. We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf. We employ same training steps for fair comparison. For MathVista, the reward function consists of format and accuracy rewards. For other tasks, we adopt the rewards from VisionReasoner [21], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
@@ 4.2 Comparison of Different Training Paradigms
Main Results. A comparative analysis of post-training paradigms across various domains is presented in Table 1. Empirical results indicate that ViSurf consistently outperforms existing post-training paradigms across various domains, with an average relative gain of 38.6% over the baseline model. This advantage is particularly significant in domains where the baseline model demonstrates lower competency, such as Non-Object and Anomaly, suggesting the method’s efficacy in addressing domain exceeding model’s knowledge base. However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal. We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during pre-training. In comparison, both RLVR and ViSurf preserve the original model performance. Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model. We attribute this to the fact that self-generated rollouts frequently produce incorrect answers, thereby hindering model optimization.
Catastrophic Forgetting. We evaluate the performance of ChartQA [25] and DocVQA [26]. As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms. Both RLVR and ViSurf demonstrate robustness against catastrophic forgetting. In contrast, SFT and SFT →RLVR suffer from performance degradation, which is attributable to catastrophic forgetting.
@@
ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B [38]. As shown in Table 3, our method consistently outperforms its counterparts across different models. Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
@@ 4.3 Ablation of Reward Control
Aligning Ground-truth Labels with Rollouts Preference. The empirical results substantiate the critical importance of this strategy, as evidenced by a consistent performance decline across multiple datasets upon its ablation. This observation provides strong empirical validation for the theoretical requirement of $\pi_{\theta}\approx\pi_{\theta_{old}}$ presented in Equation 10.
@@
Smoothing the Reward for Ground-truth Labels. The performance decline observed across datasets following the ablation of reward smoothing underscores its necessity. Concurrently, the results suggest that the SFT Term in Equation 10 becomes superfluous and can be removed when the model’s self-rollouts already achieve a higher-quality solution.
@@ 4.4 In-depth Analysis
Boundary Analysis. As shown in Table 1, the performance gain from our method is contingent on the baseline model’s initial performance. When the baseline performs poorly (e.g., below 50%), indicating its inadequacy for the task, our method yields a substantial improvement. Conversely, when the baseline already achieves high performance (e.g., above 50%), signifying a strong starting point, the upper bound of our method aligns with that of RLVR alone. This observation corroborates our theoretical analysis in Section 3.5.
@@
Reduce the Burden of Prompt Design. The RLVR paradigm relies heavily on carefully engineered user prompts to guide the model toward generating rollouts in a specific format, often requiring explicit instructions such as output with format ’point_2d’: [2, 3]. In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering. Table 5 compares performance with and without detailed prompts (appendix B), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
@@
Training Cost. We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms. Each method was implemented using well-established frameworks: DeepSpeed [28] and TRL [6] for SFT, and VeRL [33] for RLVR and ViSurf. The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.
@@
4.5 Comparison with State-of-The-Arts
We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg. We compare LISA [12], GSVA [39], SAM4MLLM [2], SegZero [20], VisionReasoner [21]. As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
@@ 5 Conclusion
6 Discussion and Future Work
@@
Supplementary Material
Appendix A Additional Explanation on pure RLVR
@@
Appendix B Illustration of Prompt Design
Table 9 shows prompt with detailed format instruction, where we provide desired output format for model. In contrast, Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.

### v2

@@ 4.1 Experimental Settings
We verify ViSurf on benchmarks across several domains.
Non-Object Segmentation. The gRefCOCO [liu2023gres] includes queries that do not contain corresponding objects. The evaluation metrics are gIoU and N-Acc. We use Multi-objects-7K plus with 200 non-object data for training.
Reasoning Segmentation. The ReasonSeg [lai2024lisa] includes test samples that need reasoning for correct segmentation. It has 200 validation images and 779 test images. The evaluation metric is gIoU. We use Multi-objects-7K proposed in VisionReasoner [liu2025visionreasoner] for training.
GUI Grounding. The OmniACT [kapoor2024omniact] is a GUI grounding task for Desktop and Web. We derive 6,101 samples in training split and verify on the test split. The accuracy is calculated as whether the predict point correctly locates inside the interest region.
Anomaly Detection. The RealIAD [wang2024real] includes real-world, multi-view industrial anomaly. We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint. We calculated the ROC_AUC.
Medical Image: Skin. The task one of ISIC2018 [codella2019skin, isic2018] is lesion segmentation. It includes 2,594 training samples and 1,000 test samples. We measure the bbox_acc metric, which computes the ratio of predicted bounding boxes whose IoU with the ground truth exceeds 0.5.
MathVista. The MathVista-testmini [lu2023mathvista] includes 1,000 diverse mathematical and visual tasks. We gather around 10k training data from WeMath [qiao2024we], MathVision [wang2024measuring], Polymath [gupta2024polymath], SceMQA [liang2024scemqa], Geometry3K [lu2021inter].
Implemention Details. We instantialize ViSurf algorithm with Qwen2.5VL-7B [bai2025qwen25vl] and adopt SAM2 [ravi2024sam2] if needed. We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf. We employ same training steps for fair comparison. For MathVista, the reward function consists of format and accuracy rewards. For other tasks, we adopt the rewards from VisionReasoner [liu2025visionreasoner], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
@@ 4.2 Comparison of Different Training Paradigms
Main Results. A comparative analysis of post-training paradigms across various domains is presented in Table 1. Empirical results indicate that ViSurf consistently outperforms existing post-training paradigms across various domains, with an average relative gain of 38.6% over the baseline model. This advantage is particularly significant in domains where the baseline model demonstrates lower competency, such as Non-Object and Anomaly, suggesting the method’s efficacy in addressing domain exceeding model’s knowledge base. However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal. We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during baseline model’s pre-training stage. In comparison, both RLVR and ViSurf preserve the baseline model’s performance. Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model. We attribute this to the fact that self-generated rollouts frequently produce incorrect answers, thereby hindering model optimization.
Catastrophic Forgetting. We evaluate the performance of ChartQA [masry2022chartqa] and DocVQA [mathew2021docvqa]. As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms. Both RLVR and ViSurf demonstrate robustness against catastrophic forgetting. In contrast, SFT and SFT → RLVR suffer from performance degradation, which is attributable to catastrophic forgetting.
@@
ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B [wang2024qwen2vl]. As shown in Table 3, our method consistently outperforms its counterparts. Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
@@ 4.3 Ablation of Reward Control
Aligning Ground-truth Labels with Rollouts Preference. The empirical results underscore the critical importance of this strategy, as its ablation leads to consistent performance degradation across multiple datasets. This observation provides strong empirical validation for the theoretical requirement of $\pi_{\theta}\approx\pi_{\theta_{old}}$ presented in Equation 10.
@@
Smoothing the Reward for Ground-truth Labels. The performance decline observed across datasets following the ablation of reward smoothing underscores its necessity. Concurrently, the results suggest that the SFT Term in Equation 10 becomes superfluous when the model’s self-rollouts already achieve a higher-quality solution.
@@ 4.4 In-depth Analysis
Boundary Analysis. As shown in Table 1, the performance gain of ViSurf is related to the baseline model’s performance. When the baseline performs poorly (e.g., below 50%), indicating its inadequacy for the task, our method yields a substantial improvement. Conversely, when the baseline already achieves high performance (e.g., above 50%), signifying a strong starting point, the upper bound of our method aligns with that of RLVR alone. This observation corroborates our theoretical analysis in Section 3.5.
@@
Reduce the Burden of Prompt Design. The RLVR paradigm relies heavily on explicit instructions to guide the model toward generating rollouts in a specific format, such as output with format ’point_2d’: [2, 3]. In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering. Table 5 compares performance with and without detailed prompts (details are provided in Appendix D of the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
@@
Training Cost. We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms. Each method was implemented using well-established frameworks: DeepSpeed [deepspeed] and TRL [trl] for SFT, and VeRL [verl] for RLVR and ViSurf. The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.
4.5 Comparison with State-of-The-Arts
We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg. We compare LISA [lai2024lisa], GSVA [xia2024gsva], SAM4MLLM [chen2024sam4mllm], SegZero [liu2025seg], VisionReasoner [liu2025visionreasoner]. As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
@@ 4.5 Comparison with State-of-The-Arts
Figure 7: Visualization of ViSurf on various tasks.
4.6 Qualitative Results
Qualitative results on various tasks are presented in Figure 7. The proposed method, ViSurf, successfully localizes normal objects, identifies anomalies, detects non-object cases, and solves mathematical problems.
@@ 5 Conclusion
[…]
Supplementary Material
Appendix A Discussion and Future Work
@@
Appendix B Additional Explanation on pure RLVR
@@
Appendix C Non Object Training Data
The training data for gRefCOCO was adapted from the VisionReasoner training data [liu2025visionreasoner], from which we utilized 7k referring expression samples. To enhance the model’s ability to handle cases where no target object is present, we augmented this dataset with 200 non-object examples. These negative samples were generated by providing a question that is unanswerable given the image content and training the model to output an empty list (<answer>[]</answer>).
Appendix D Illustration of Prompt Design
Table 9 shows prompt with detailed format instruction, where we provide desired output format for model. Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.